A fountain in a shopping mall has two parabolic arcs of water intersecting. The equation of one parabola is y=-0.25x^2+2x and the equation of the second parabola is y=-0.25x^2+4x-11.75. How high above the base of the fountain do the parabolas intersect: All dimensions are in feet. Round to the nearest hundredth.
Answer by lwsshak3(11628)

A fountain in a shopping mall has two parabolic arcs of water intersecting. The equation of one parabola is y=-0.25x^2+2x and the equation of the second parabola is y=-0.25x^2+4x-11.75. How high above the base of the fountain do the parabolas intersect: All dimensions are in feet. Round to the nearest hundredth.
**
When parabolas intersect, the equations are equal to each other.
-0.25x^2+2x=-0.25x^2+4x-11.75
2x=11.75
x=11.75/2=5.875
y=-.25*(5.875)^2+2*5.875=3.121 ft
ans:
parabolas intersect 3.121 ft above the base